OpenStudy (ashley1nonly):

find the value of the derivative for the given function dr/dtheta if r= 1/ sqrt(11-theta)

OpenStudy (freckles):

you can use the quotient rule

OpenStudy (freckles):

or if you want write r as \[r=(11-\theta)^\frac{-1}{2}\] then just use chain rule

OpenStudy (ashley1nonly):

so i just use the trig functions

OpenStudy (ashley1nonly):

theta=2

OpenStudy (ashley1nonly):

how would i find a value

OpenStudy (freckles):

I don't know what you mean

OpenStudy (freckles):

you don't need trig functions to differentiate the r equation above

OpenStudy (freckles):

just chain rule

OpenStudy (ashley1nonly):

find the value of the derivative is it asking for a number or just find the derivative

OpenStudy (freckles):

find the value of the derivative for what value of theta? theta=2? you need to first find the derivative?

OpenStudy (ashley1nonly):

1/54 is final anser 1/2(11-x)^(3/2)

OpenStudy (freckles):

\[\frac{dr}{d \theta}=\frac{1}{2 (11-x)^\frac{3}{2}} \\ \frac{dr }{ d \theta} |_{x=2}=\frac{1}{2(11-2)^\frac{3}{2}} \]

OpenStudy (freckles):

so you said 11-2=9 and sqrt(9)=3 and 3^3=27 and 2(27)=54 so we have 1/54 food job

OpenStudy (freckles):

good job*

OpenStudy (ashley1nonly):

thankss